OpenStudy (errinss):

what is 1.21 repeating as a fraction?

OpenStudy (errinss):

@jigglypuff314 can you help me?

TheSmartOne (thesmartone):

Which part is repeating?

OpenStudy (errinss):

.21

TheSmartOne (thesmartone):

Here's an example on how to do it: \(x = 0.\bar{5}\) (where 5 is repeating) Multiply both sides by 10 \(10x = 5.\bar{5}\) Now, subtract \(0.\bar{5}\) from both sides \( 10x - 0.\bar{5} = 5.\bar{5} - 0.\bar{5}\) \( 10x - 0.\bar{5} = 5\) Remember in the beginning we said that \( x = 0.\bar{5}\), replace it on the left hand side \(10x - x = 5\) \( 9x = 5\) \( x = \frac{5}{9}\) So \(0.\bar{5}\) is equal to \(\Large\frac{5}{9}\) Use this method to change \( 0.\bar{21}\) into a fraction and then add 1 to the fraction since we were given \(1.\bar{21}\)